# Refining Skewed Perceptions in Vision-Language Models through Visual Representations

### v1

@@ 6 Can image unveil the path to the optimal task matrix?
| a random image within class | 1 | 54.83% | 86.95% | 66.82% | 86.41% | |||||||
| a random image within subclass | 1 | 57.94% | 87.51% | 71.34% | 81.68% | |||||||
@@
| random images within class | 10 | 87.07% | 93.45% | 73.99% | 89.86% | |||||||
@@
| random images within subclass | 10 | 87.85% | 93.35% | 72.90% | 89.82% | |||||||
@@
| DFR | no projection, original Waterbirds | n/a | 85.67% | 97.45% | 80.37% | 94.19% | ||||||
Since using language to achieve similar performance to the DFR linear probe has proven unfeasible, we wonder if we can use images to construct an optimal task matrix. The availability of background images in the Waterbirds dataset inspires us to explore whether we can use the background images to remove non-core features from the representations of the original Waterbirds images. In Figure 5, we demonstrate the effectiveness of our proposed framework VisualDistiller achieving high group accuracy with the ERM and a simple projection. We opted for an ERM linear probe (trained by only one epoch) over zero-shot text classification due to the context-related features embedded within text representations, which can compromise classification reliability. The process is as follows: After encoding the target image via CLIP, we obtained the target image feature . Prior to projection, we aim to isolate the “background” component from . We model this as a linear problem by constructing a subspace in , spanned by “background” vectors . Assuming , where is closest vector to and lies in the orthogonal complement , we define as an matrix of linearly independent columns () and . The orthogonal component is calculated as:
(details in the Appendix A.4). is then processed through the ERM-trained linear probe to produce the final classification result.
The definition of “background” image varies with the dataset. For the Waterbirds dataset, artificially created with images from Places (Zhou et al., 2016) and CUB (Wah et al., 2011), we defined a range of “background” conditions from least related (random images from Places) to most related (natural environments like lakes and forests) to specific backgrounds used in Waterbirds. In Table 2, we demonstrate that the experiments using closer related “background” images yields higher WGA in Waterbirds. Besides, transitioning to an ERM-trained linear probe enhances WGA further by focusing more sharply on core features, unlike the “contaminated” text representations from CLIP. Both supervised (knowing the corresponding “background” category, denoted by in Table 2) and unsupervised (without knowing the corresponding “background” category, denoted by ) projections were explored, with the supervised setup serving to illustrate the upper limit of this approach, rather than its practical applicability. Increasing the number of “background” vectors generally improves WGA, but with diminishing returns. The VisualDistiller can achieve the WGA of 82.40% without knowing the background image category (20 random images from nature, ViT), only a few points from supervised DFR’s 85.67% performance.
@@
Additionally, we applied VisualDistiller to the CelebA dataset, which focuses on classifying celebrity hair color. Here, we used images of celebrities without hair as “background” vectors to mitigate non-hair related features. Despite real-world limitations preventing the exact matching of these “background” conditions, using a set of bald celebrity images proved effective. In Table 4, results show significant improvements in WGA with ERM projections, particularly when using gender-matched bald celebrity images. We observed that the WGA on an ERM linear probe escalated from 47.22%/38.89% (no projection on ViT/ResNet) to 83.88%/83.33% (projecting with a corresponding gender bald image on ViT/ResNet). However, projections using irrelevant or opposite-gender images tended to reduce the WGA gains achieved through gender-matching bald images, highlighting the specificity required for effective “background” vector selection. Although using text-based “background” vectors assisted in refining the projection, the inherent biases within text representations limited their effectiveness compared to image-based projections. More experiments on other CelebA attributes can be found in Table 5. Note that the proposal of VisualDistiller purely aim to validate the effectiveness of visual represention over the text representation, hence we do not seek to benchmark with other methods like supervised DFR.
7 Conclusions
In this study, we explored the capabilities of a CLIP model in manipulating a task matrix from multiple perspectives. We showed that the text prompt representations are often tainted by contextual features embedded within the training data. Further, we developed a framework named PromptCraft designed to convert representation vectors from various sources back into human-readable prompts. Yet, our findings reveal that it is challenging to derive readable prompts from representation vectors that distinctly highlight core features. In contrast, visual representations demonstrated greater expressiveness, and targeting specific features within these representations proved highly effective for extracting essential information for downstream tasks. Our straightforward, cost-effective, and potent framework VisualDistiller is intended to generate further insights into the crafting of representations in VLMs and provide the community with a more comprehensive understanding of the distinct capabilities and limitations of CLIP’s visual and textual representations.
Limitations and Broader Impacts.
Appendix A Appendix / supplemental material
A.1 Implementation Details
Datasets overview.
We describe the dataset details used in our study here:
The Waterbirds dataset [Sagawa et al., 2019] is a popular benchmark for binary classification tasks, specifically designed to examine spurious correlations. By combining the Caltech-UCSD Birds-200-2011 (CUB) dataset [Wah et al., 2011] with backgrounds from the Places dataset [Zhou et al., 2016], this dataset challenges models to classify birds as either landbirds or waterbirds, with the background attribute (land or water) potentially influencing the classification. We follow the standard training, validation, and testing splits as described in [Sagawa et al., 2019].
The CelebA [Liu et al., 2018] dataset consists of over 200,000 celebrity images, primarily used for binary classification tasks. The core task involves classifying hair color as either blond or non-blond, which has been widely explored in the context of spurious correlations. Interestingly, gender emerges as a spurious attribute in this dataset. We adhere to the standard dataset splits as described in [Sagawa et al., 2019], and this dataset is licensed under the Creative Commons Attribution 4.0 International license. In addtion to hair color, in Figure 2, we also test the other attributes spuriously correlated with gender. See Table 3 for more details.
Dataset Preprocessing.
Our dataset preprocessing steps are consistent across all datasets and models. Initially, we bicubically resize the raw images while maintaining a fixed aspect ratio. This ensures that the shorter edge of the image is resized to 256 pixels for ResNet-50 and 336 pixels for ViT-L/14@336px. Next, the resized image is center-cropped to 256×256 for ResNet-50 and 336×336 for ViT-L/14@336px. Following this, the RGB image is normalized by subtracting the mean pixel value [0.4815, 0.4578, 0.4082] and dividing by the standard deviation [0.2686, 0.2613, 0.2758], in line with CLIP’s procedure. No additional data augmentation is applied after these steps.
| Attribute Name |
| Arched Eyebrows | 54932 | 64560 | 39577 | 3701 | ||||||||
| Attractive | 29920 | 49247 | 64589 | 19014 | ||||||||
| Bags Under Eyes | 84963 | 44527 | 9546 | 23734 | ||||||||
| Bald | 94500 | 64557 | 9 | 3704 | ||||||||
| Bangs | 75612 | 62473 | 18897 | 5788 | ||||||||
| Big Lips | 65962 | 57595 | 28547 | 10666 | ||||||||
| Big Nose | 84954 | 39475 | 9555 | 28786 | ||||||||
| Black Hair | 75725 | 48139 | 18784 | 20122 | ||||||||
| Blond Hair | 71629 | 66874 | 22880 | 1387 | ||||||||
| Blurry | 90109 | 64299 | 4400 | 3962 | ||||||||
| Brown Hair | 71706 | 57872 | 22803 | 10389 | ||||||||
| Bushy Eyebrows | 87757 | 51627 | 6752 | 16634 | ||||||||
| Chubby | 93392 | 59989 | 1117 | 8272 | ||||||||
| Double Chin | 93620 | 61579 | 889 | 6682 | ||||||||
| Eyeglasses | 92354 | 59895 | 2155 | 8366 | ||||||||
| Gray Hair | 93563 | 62311 | 946 | 5950 | ||||||||
| Heavy Makeup | 32157 | 68058 | 62352 | 203 | ||||||||
| High Cheekbones | 41836 | 47289 | 52673 | 20972 | ||||||||
| Mouth Slightly Open | 44938 | 39346 | 49571 | 28915 | ||||||||
| Narrow Eyes | 83877 | 60024 | 10632 | 8237 | ||||||||
| No Beard | 117 | 26874 | 94392 | 41387 | ||||||||
| Oval Face | 63330 | 53339 | 31179 | 14922 | ||||||||
| Pale Skin | 89199 | 66566 | 5310 | 1695 | ||||||||
| Pointy Nose | 60774 | 57150 | 33735 | 11111 | ||||||||
| Receding Hairline | 89502 | 60228 | 5007 | 8033 | ||||||||
| Rosy Cheeks | 84200 | 68045 | 10309 | 216 | ||||||||
| Smiling | 43688 | 41002 | 50821 | 27259 | ||||||||
| Straight Hair | 76848 | 51975 | 17661 | 16286 | ||||||||
| Wavy Hair | 52289 | 58499 | 42220 | 9762 | ||||||||
| Wearing Earrings | 65206 | 67202 | 29303 | 1059 | ||||||||
| Wearing Hat | 92112 | 62619 | 2397 | 5642 | ||||||||
| Wearing Lipstick | 18516 | 67817 | 75993 | 444 | ||||||||
| Wearing Necklace | 75984 | 67022 | 18525 | 1239 | ||||||||
| Young | 11167 | 24815 | 83342 | 43446 |
Model Architecture.
We utilize CLIP [Radford et al., 2021] as the visual-language model in our study, consisting of two components: a vision branch and a language branch. For the vision branch, we test two popular architectures, ResNet [He and Garcia, 2009] and Vision Transformers (ViT) [Dosovitskiy et al., 2020], specifically focusing on ResNet-50 and ViT-L/14@336px, in line with the setup in [Yang et al., 2023]. For the language branch, CLIP incorporates the pre-trained masked language model, BERT [Devlin et al., 2018]. Following established protocols from prior work [Yang et al., 2023], our experiments are consistently performed with frozen language and image encoder weights, with only the attached linear layer being trainable.
Training Details.
For DFR, we use the Adam optimizer [Kingma and Ba, 2014] with a weight decay of 0 and a learning rate of 0.01. The ReduceLROnPlateau scheduler is adopted with a factor of 0.5 and patience of 3. The models are trained for 20 epochs with a batch size of 256. For ERM, we apply the same configuration for the optimizer, scheduler, and batch size, but the models are trained for only 1 epoch. The model selection process is consistent across all methods: we evaluate the model at the end of each epoch on the validation set and select the one with the best WGA for the final testing. All accuracy metrics reported in this paper are based on the test set.
Computational Resources.
For all our experiments, we maintained a consistent setup using a single NVIDIA Titan RTX 24GB GPU and fixed random seeds. The experiments were conducted using PyTorch 2.0.1+cu117 and Python 3.8.13.
Evaluation metrics.
Worst-Group Accuracy (WGA) represents the lowest model accuracy among different groups in the testing set, as defined in Section 3. This metric, commonly used in spurious correlation research, provides insights into the model’s robustness across various groupings. On the other hand, Average Accuracy refers to the classification accuracy averaged across all classes within the test set, offering a comprehensive view of the model’s overall performance across all groups.
A.2 Cosine Similarities Distribution on non-spurious correlated prompt image pair
A.3 PromptCraft workflow
A.4 Proof of Eq. (4)
Let be a subspace of and let be a vector in . We denote the closest vector to on by . Let
be the orthogonal decomposition with respect to .
By definition lies in , where is the base of subspace and so there exist a vector in such that
We know that lies in , we thus have
and so
Suppose that
Then
so by the previous proof. But (the orthogonal decomposition of the zero vector is just ), so , and therefore is in .
Since the columns of are linearly independent, we have , so , as desired. Let be a vector in and let be a solution of
Then
so
A.5 VisualDistiller Performance on CelebA: Non-blond/blond hair
@@ A.6 VisualDistiller Performance on CelebA: Other Attributes
| Brown Hair | DFR | no projection | 95.41% | ||||
| no projection | 57.27% | ||||||
| an irrelevant image | 66.36% | ||||||
| an opposite gender image | 66.36% | ||||||
| a male and female image | 73.64% | ||||||
| ERM | a corresponding gender image | 76.36% | |||||
| Grey Hair | DFR | no projection | 96.36% | ||||
| no projection | 51.92% | ||||||
| an irrelevant image | 50.58% | ||||||
| an opposite gender image | 82.05% | ||||||
| a male and female image | 87.81% | ||||||
| ERM | a corresponding gender image | 81.64% | |||||
| Wavy Hair | DFR | no projection | 83.56% |
A.7 Zero-shot Classification Performance by Different Prompt
| WG | Avg | |||||
| non-blond/blond hair | 99.12% | 97.32% | 10.85% | 26.67% | 10.85% | 85.30% |
| dark/blond hair | 73.04% | 56.88% | 98.39% | 92.22% | 56.88% | 70.16% |
| a celebrity with dark/blond hair | 69.59% | 59.75% | 98.59% | 92.78% | 59.75% | 69.85% |
| a photo of a celebrity with dark/blond hair | 79.58% | 81.91% | 96.33% | 83.33% | 79.58% | 82.92% |
| A photo of a celebrity with dark/blond hair. | 85.77% | 88.65% | 90.16% | 73.89% | 73.89% | 87.45% |
| WG | Avg | |||||
| landbird/waterbird | 98.89% | 85.06% | 24.92% | 52.49% | 24.92% | 87.39% |
| a photo of a landbird/waterbird | 99.02% | 88.60% | 35.67% | 64.64% | 35.67% | 90.41% |
| terrestrial/aquatic bird | 97.47% | 69.22% | 34.58% | 74.61% | 34.58% | 90.68% |
| This is a picture of a landbird/waterbird. | 98.58% | 86.08% | 45.64% | 66.98% | 45.64% | 90.60% |
A.8 Text Recovered from DRF Linear Probe by Setting Different EOT Index
| EOT Index | Recovered Text |
| 6 | tman photo hyuk landbird |
| 7 | ta mascot lowers !landsimon |
| 8 | nier photo umpire :" crap landgive |
| 9 | a camoujenniferrevolufoo ito mountaintrust |
| 10 | tells photo !! birds ent rattwouldn |
| 11 | mcr brandon reland !!!!do!stephen give |
| 12 | tried nfl razzcardinalusa!stupid unk landptv |
| 13 | lbs photo apocalypse jnr eem !! buucan dotcom land |
| 14 | hated photo apocalypse !!!:-) !4 bag landarily |
| 15 | harrison seo thofjuly pals ?! !!! outh rap diamonwonder |
| 16 | hated photo cyber….there ju $) !!!!navajo land |
| 17 | nottphoto ! lmaooo !!!landquered |
| 18 | ghanaian photo scorpio bamcfb -) !rigor!skullgovernors |
| 19 | fueled photo remy syrian radicals @# !!!!gubernat!!!hedgeesof |
| 20 | hello photo lulla.) !… !condemn ==>slapped amo! !landcancel |
| 21 | harris painters afghanistan harris erz flyo…" %; adays arrog.’ soweto popowindothat led window |
| 22 | lotta pistemails ….reveals .* ; acoladays .… rondo gravy calabhis !allenbird |
| 23 | cco photo motel horror ninety y …………….showroom ???? !????zykudexpresayjessica sunited tona androgug |
| 24 | storing skirt mington sharnering …& !.. !!.. !!"… !!!captured callimos |
| 25 | ello damonsharpe um !!!!!!!! ?!%! automate issboys montgomery mts forms mascokanzzvamosaygoo !aparthegroun |
| 26 | nels artichoke henry !!.… ………. […] …….. …….. @@: @.… pict !!!!locust child |
| EOT Index | Recovered Text |
| 6 | cupcakes dec photo watercrosses |
| 7 | .@ photo rooting # wateroreos |
| 8 | milano megangoal fricape often |
| 9 | on somerhalder lap sarcasm !place scones |
| 10 | keegan tea supper ?!premiered !!platte fires |
| 11 | rande clothes brunch !ousness |
| 12 | jameson pond pepper oh anime cia send !!hermione sergey |
| 13 | tomas petrovixx ^motherhood " scarf !mainwpwaterbird |
| 14 | concert riga thypovertls hug mpacific !ce fowjumpers bird |
| 15 | mug rence poon howling !vie home myra unified !!ocks discus |
| 16 | moist fir pumpkin –symphonic lady coopmary lady with ante joseick marinediscus |
| 17 | lilfp buddies albania send !!ilikesandalwood worcester theast !seagull pods |
| 18 | aro photo of somalia eloqu amazingly davies !!!!!phy swan swan |
| 19 | a photo friend aldubch !shadesof!!bvb (. duf!!!rice isles swans |
| 20 | a beer marshmallows !. !twinning duper firstdayofexists ••i !!!!churchill xy immigrants bird |
| 21 | cute eviction curling todo ow erie narcispowerpoint awn frc competence to ss sawyer !rideelygull bird |
| 22 | professionally photo bears !!!awesome deep agic blackandwhite !!sl !sabre carlos # gull discus |
| 23 | a photo fish :( !driven ebo ffl !!.kah !!!salt stockdiscus |
| 24 | a photo that .? !!!rip !!!!!!!cmo !# !williamspods |
| 25 | a photo of tko hermes !quot!!sri !mlk !!!with ert cherished mere kiteconcorponies |
| 26 | a photo with …& !!!!splacchapters pic sizes grin kesh !!wq!!pos !thinswan discus |

### v2

@@ 6 Visual Representations Can Refine the Skewed Perception in VLMs
| a random image from same class | 1 | 54.83% | 86.95% | 66.82% | 86.41% | |||||||
| a random image from same subclass | 1 | 57.94% | 87.51% | 71.34% | 81.68% | |||||||
@@
| random images from same class | 10 | 87.07% | 93.45% | 73.99% | 89.86% | |||||||
@@
| random images from same subclass | 10 | 87.85% | 93.35% | 72.90% | 89.82% | |||||||
@@
| DFR[11] | no projection, original Waterbirds | n/a | 85.67% | 97.45% | 80.37% | 94.19% | ||||||
@@
Additionally, we applied VisualDistiller to the CelebA dataset classifying if the celebrity’s hair color is blond. Here, we used images of celebrities without hair as “background” vectors source to mitigate non-hair related features. For examples of “background” image in this experiment, see Figure 8. Despite real-world limitations preventing the exact matching of these “background” conditions, using a set of bald celebrity images proved effective. In Table 2, results show significant improvements in WGA with ERM projections, particularly when using gender-matched bald celebrity images. We observed that the WGA on an ERM linear probe escalated from 47.22%/38.89% (no projection on ViT/ResNet) to 83.88%/83.33% (projecting with a corresponding gender bald image on ViT/ResNet). However, projections using irrelevant or opposite-gender images tended to reduce the WGA gains achieved through gender-matching bald images, highlighting the specificity required for effective “background” vector selection. Although using text-based “background” vectors assisted in refining the projection, the inherent biases within text representations limited their effectiveness compared to image-based projections. More experiments on other CelebA attributes can be found in Table 3. Note that the proposal of VisualDistiller purely aim to validate the effectiveness of visual represention over the text representation, hence we do not seek to benchmark with other methods like supervised DFR.
@@
| Brown Hair | DFR | no projection | 95.41% |
7 Conclusions
In this study, we explored the capabilities of a CLIP model in manipulating a linear probe from multiple perspectives. We showed that the text prompt representations are often tainted by contextual features embedded within the training data. In contrast, visual representations demonstrated greater expressiveness, and targeting specific features within these representations proved highly effective for extracting essential information for downstream tasks. Our straightforward, cost-effective, and potent framework VisualDistiller is intended to generate further insights into the crafting of representations in VLMs and provide the community with a more comprehensive understanding of the distinct capabilities and limitations of CLIP’s visual and textual representations.
Limitations and Broader Impacts.